# Personalized Item Representations in Federated Multimodal Recommendation

### v2

@@ Abstract.
Federated recommendation systems are essential for providing personalized recommendations while protecting user privacy. However, current methods mainly rely on ID-based item embeddings, neglecting the rich multimodal information of items. To address this, we propose a Federated Multimodal Recommendation System, called FedMR. FedMR uses a foundation model on the server to encode multimodal item data, such as images and text. To handle data heterogeneity caused by user preference differences, FedMR introduces a Mixing Feature Fusion Module on each client, which adjusts fusion strategy weights based on user interaction history to generate personalized item representations that capture users’ fine-grained preferences. FedMR is compatible with existing ID-based federated recommendation systems, improving performance without modifying the original framework. Experiments on four real-world multimodal datasets demonstrate FedMR’s effectiveness. The code is available at https://anonymous.4open.science/r/FedMR.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Collaborative filteringccs: Information systems Personalizationccs: Information systems Combination, fusion and federated search
@@ 1. Introduction
Under privacy protection regulations, such as GDPR (Voigt and Von dem Bussche, 2017), that emphasize user data privacy, Federated Recommendation Systems (FedRec) have become an essential technology in modern life (Sun et al., 2024). However, in these systems, since each client typically only has access to a single user’s interaction history, data sparsity is a common issue that degrades recommendation performance (Li and Long, 2024). Modeling item features as shared information across clients can effectively mitigate the impact of data sparsity. Many studies (Zhang et al., 2023c; Li et al., 2023a, 2024b) have shown that personalized item representations can significantly improve recommendation results. Despite items often containing rich multimodal information (Zhou et al., 2023b), such as such as text and images, most existing FedRec methods (Zhang et al., 2023d, 2024d, 2024f) still rely mainly on ID embeddings, underutilizing the potential of multimodal data. Thus, generating personalized item representations from multimodal data while ensuring user privacy remains a significant challenge.
As a potential solution, Multimodal Recommendation Systems (MMRec) integrate data from different modalities, significantly enhancing recommendation performance (He and McAuley, 2016; Wei et al., 2019; Zhou et al., 2023a). Unlike single-modal systems, MMRec can capture hidden relationships and complementary information between various modalities, leading to a better understanding of user preferences and item characteristics (Zhou et al., 2023b). For example, MMRec methods (Wei et al., 2019; Wang et al., 2021; Tao et al., 2022) combine video frames, audio tracks, and item descriptions when users interact with online platforms, providing a more accurate representation of user preferences. To effectively handle multimodal data, MMRec often uses Foundation Models (FM) pretrained on large-scale datasets, which possess powerful feature extraction and representation capabilities (Bommasani et al., 2021). By applying FMs, MMRec can efficiently extract and integrate information from multiple modalities, capturing commonalities and complementary relationships between them (Radford et al., 2021a; Cui et al., 2022; Geng et al., 2022, 2023; Wu et al., 2024). However, these centralized MMRec methods rely on servers to aggregate user data, digital activities, and preferences to train accurate models (Liu et al., 2024), which raises privacy concerns (Chai et al., 2020; Li et al., 2023a).
To address these challenges, Federated Multimodal Recommendation Systems (FedMMRec) have emerged, which integrates MMRec and introduces the capability of processing multimodal information into FedRec. By utilizing multimodal data and user-specific information, such as profiles and interaction history, locally on each client to generate personalized recommendations and aggregating model parameters at the server, FedMMRec enhances recommendation performance while preserving user privacy (Feng et al., 2024). To the best of our knowledge, FedMMRec is still in the early stages of research, and related studies are limited. FedMMR (Wu et al., 2024) adopts a model-splitting strategy where the computationally intensive components are handled by the server, while clients perform lightweight collaborative filtering. AMMFRS (Feng et al., 2024) introduces an attention mechanism on the client side to perform weighted fusion of multimodal data, effectively distinguishing and utilizing the contributions of different modalities. These methods not only protect user privacy but also effectively integrate multimodal features to improve recommendation accuracy and personalization. However, they rely on fixed multimodal fusion strategies, neglecting the personalized representation of item features on clients, which limits their ability to adapt to diverse user needs. Although AMMFRS uses an attention mechanism to distinguish the importance of different modalities, it lacks adaptability to the diversity of user data, resulting in limited flexibility and accuracy of fusion strategies (Niu et al., 2021).
Main Contribution. To integrate multimodal item information into FedRec while ensuring user privacy, we propose a novel Federated Multimodal Recommendation system (FedMR). FedMR uses a foundation model on the server to encode multimodal item data (e.g., images, text) into feature embeddings, and constructs a global item ID embedding to capture structural features among items. On each client, FedMR introduces a Mixing Feature Fusion Module that dynamically adjusts the importances of fusion strategies based on user interaction history, combining multimodal and ID embeddings to model fine-grained personalized item representations for users. Thus, FedMR seamlessly integrates multimodal information with existing ID-based FedRec methods, improving the flexibility and accuracy of recommendations without altering the existing framework, and effectively handles data heterogeneity across clients, providing more personalized recommendations.
The main contributions of this work are summarized as follows:
  • We propose FedMR, a federated multimodal recommendation system using foundation models, which efficiently leverages multimodal data while preserving user privacy;
  • A Mixing Feature Fusion Module is designed to dynamically adjusts fusion strategies based on user interaction histories to generate personalized item representations, enhancing personalization and flexibility in recommendations;
  • FedMR is compatible with existing ID-based federated recommendation methods, improving recommendation performance without modifying the existing framework;
  • Extensive experiments are conducted on four real-world multimodal recommendation datasets to validate the effectiveness of our proposed method FedMR;
  • We provide the source code to support the reproducibility.
@@ 2.1. Federated Recommendation Systems
Federated recommendation systems (FedRec) aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020). In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024a), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d). However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely solely on ID embeddings to model item representations, overlooking the rich multimodal information (e.g., titles and covers) that items may have. AMMFRS (Feng et al., 2024) and FedMMR (Wu et al., 2024) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec. AMMFRS uses an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data, emphasizing the contribution of different features. In contrast, FedMMR assigns the computationally heavy multimodal learning tasks to the server and introduces a secure cross-client convolution strategy to capture high-order interaction information using secret sharing techniques while preserving privacy. Unlike methods that focus only on personalizing user information, PFedRec (Zhang et al., 2023c) proposes a dual personalization mechanism that optimizes personalization for both users and items, integrating it into existing FedRec frameworks. Building on this, FedRAP (Li et al., 2023a) introduces additive personalization, which refines the modeling of user and item features. FedDAE (Li et al., 2024b) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network. While these methods achieve personalized item representations, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
@@ 2.2. Multimodal Recommendation Systems
2.3. Foundation Models
@@ 3. Problem Formulation
Let $\mathcal{U}=\{u\}$ and =$\mathcal{I}=\{i\}$ be the set of users and items correspondingly, we assume that the users’ interaction history data is implicit feedback, represented by $\mathbf{R}=[\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{n}]^{T}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|},$ where $r_{u,i}=1$ if user $u$ has interacted with item $i$, and $r_{u,i}=0$ otherwise. For user $u$, we use $\mathbf{r}_{u}\in\{0,1\}^{|\mathcal{I}|}$ to denote their interaction history and $\mathcal{I}_{u}$ to represent the set of items they have interacted with. Moreover, we define $\mathcal{O}=\{(u,i)|\text{user }u\text{ has interacted with item }i\}$ to index the items each user has interacted with. Each item in $\mathcal{I}$ has a set of modality features $\mathcal{M}$. The $m$-th modality feature is represented as $\mathbf{E}_{m}=[\mathbf{e}_{1,m},\mathbf{e}_{2,m},\dots,\mathbf{e}_{|\mathcal{I}|,m}]$. To capture user diversity, FedMR also introduces a set of modality fusion strategies $\mathcal{G}=\{g_{1},g_{2},\cdots,g_{|\mathcal{G}|}\}$. In this work, we focus on the visual and textual modalities, denoted as =$\mathcal{M}=\{v,c\}$, respectively. It is important to note that FedMR is not limited to these two modalities and the existing fusion strategies, it can include additional modalities or modify the fusion strategies as needed.
3.1. Objective of FedRec
In classic FedRec with implicit feedback, each client only has the user’s historical interaction data $\mathbf{r}_{u}$. The goal is to train a personalized recommendation model for each user while protecting user privacy, providing personalized recommendations by predicting possible scores $\hat{\mathbf{r}}_{u}$ for items the user has not yet interacted with, based on their observed historical data $\mathbf{r}_{u}$. To measure the difference between the prediction $\hat{\mathbf{r}}_{u}$ and the truth $\mathbf{r}_{u}$, we have the following:
(1) $\min\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(\hat{r}_{u,i},r_{u,i}).$
Here, the function $\mathcal{L}_{\text{recon}}$ measures the reconstruction error between $\hat{\mathbf{r}}_{u}$ and $\mathbf{r}_{u}$, and $\alpha_{u}$ represents the loss weight for the client $u$ ($u\in\mathcal{U}$), which is used to balance the client’s contribution to the overall training, and satisfies the condition $\sum_{u\in\mathcal{U}}\alpha_{u}=1$. By optimizing Eq. (1), FedRec aims to make the predicted item scores $\hat{\mathbf{r}}_{u}$ for each user $u$ as close as possible to their true interactions ${\mathbf{r}}_{u}$.
Figure 1. The framework of FedMR. FedMR utilizes a FM to process multimodal data, including visual and textual information, to enhance personalized recommendations. On the server side, FedMR encodes visual and textual data into fixed embeddings. On each client, a Mixing Feature Fusion Module dynamically combines these embeddings with ID embeddings, adjusting the fusion weights based on user interactions to generate personalized item embeddings, capturing fine-grained user diversity.
3.2. ID-based FedRec
Currently, most FedRec methods (Lin et al., 2020a, b; Liang et al., 2021; Perifanis and Efraimidis, 2022; Zhang et al., 2023c, d; Li et al., 2023a, 2024b) construct a global item embedding $\mathbf{D}$ to capture the items’ features and their relationships with neighboring items, forming $\hat{\mathbf{r}}_{u}$ as follows:
(2) $\hat{\mathbf{r}}_{u}=f(\mathbf{D};\theta_{u}),$
where the function $f$, driven by the user-specific parameters $\theta_{u}$, generates personalized predictions $\hat{\mathbf{r}}_{u}$ upon receiving $\mathbf{D}$. However, these methods overlook the rich item multimodal information. Therefore, we consider such methods as ID-based FedRec in this work.
@@ 4. Methodology
The aim of our proposed method FedMR is to integrate the multimodal features of items with ID embeddings within a federated environment, and incorporates these features into existing ID-based FedRec methods. By doing so, we transform current FedRec methods into FedMMR methods, allowing them to fully utilize the multimodal features of items to predict items that users have not yet viewed but may find interesting, while protecting user privacy.
4.1. Encoding Multimodal Information with FM
In RecSys, users typically interact with only a small subset of all available items, which are usually stored on the server. Therefore, in multimodal scenarios, we assume the server holds the complete modal information of the item set. FedMR begins by introducing a foundation model $h$ on the server side to encode the items’ modal information, producing feature embeddings for each modality:

### v3

# Dynamic Fusion Strategies for Federated Multimodal Recommendations

@@ Abstract.
Delivering deeply personalized recommendations necessitates understanding user interactions with diverse multimedia features, but achieving this within the constraints of Federated Recommendation Systems (FedRec) is severely hampered by communication bottlenecks, user heterogeneity, and the complexity of privacy-preserving multimodal fusion. To this end, we propose FedMR, a novel multimodal FedRec framework centered around the Mixing Feature Fusion Module (MFFM). FedMR employs a two-stage process: (1) Server-side centralized multimedia content processing provides rich, shared item context using pre-trained models, mitigating limitations from client sparsity and resource constraints efficiently. (2) Client-Side Personalized Refinement, where the MFFM dynamically adapts these server-provided multimodal representations based on client-specific interaction patterns, effectively tailoring recommendations and resolving heterogeneity in user preferences towards different modalities. Extensive experiments validate that FedMR seamlessly enhances existing ID-based FedRecs, effectively transforming them into high-performing federated multimodal systems.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Collaborative filteringccs: Information systems Personalizationccs: Information systems Combination, fusion and federated search
@@ 1. Introduction
Recommendation Systems (RecSys) are integral components of modern digital platforms, guiding users through vast item catalogs by suggesting relevant content (Li and Long, 2024). Traditional RecSyss rely heavily on collaborative filtering signals derived from user-item interaction histories (Chen et al., 2017). However, the rich information embedded within item multimedia content, such as product images, video trailers, textual descriptions, and audio samples, offers a powerful complementary signal (Deldjoo et al., 2020; Zhou et al., 2023b). Leveraging these multimodal cues allows Multimodal Recommendation Systems (MMRec) (He and McAuley, 2016; Wei et al., 2019; Zhou et al., 2023a; Yu et al., 2023) to gain a deeper understanding of item characteristics and user preferences, capture subtle nuances missed by interaction data alone, significantly alleviate the cold-start problem for new items (Pan et al., 2022), and ultimately enhance recommendation accuracy and user experience.
Despite their effectiveness, centralized MMRec approaches typically require aggregating vast amounts of user interaction data and potentially sensitive item features on central servers (Liu et al., 2024). This practice raises significant privacy concerns, particularly with the enforcement of stringent data protection regulations like the GDPR (Voigt and Von dem Bussche, 2017). To address these concerns, Federated Learning (FL) (McMahan et al., 2017) has emerged as a compelling paradigm, enabling collaborative model training across distributed clients, e.g., user devices, without exposing raw local data. Federated Recommendation Systems (FedRec) (Yang et al., 2020; Wang et al., 2022) adapt FL principles to the recommendation domain, offering a promising path towards privacy-preserving personalization.
However, integrating rich multimedia information into the FL setting poses distinct challenges beyond traditional ID-based FedRecs (Feng et al., 2024). First, processing raw, high-dimensional multimedia data on resource-constrained client devices is often computationally prohibitive, hindering feature extraction and user participation (Li et al., 2024a). Second, transmitting large multimedia features incurs significant communication overhead and raises privacy concerns, as shared features might reveal sensitive information (Chai et al., 2020; Li et al., 2023a). Third, data heterogeneity (Yuan et al., 2024; Li et al., 2019) extends to user preferences for different modalities, such as visual vs. textual (Liu et al., 2019; Niu et al., 2021). This makes a single, global fusion strategy suboptimal for personalized multimodal understanding (Feng et al., 2024; Li et al., 2024a). Finally, client-side data sparsity (Sun et al., 2024) limits item understanding, especially for cold-start items (Zhang et al., 2024c). While multimedia data could alleviate this (He and McAuley, 2016; Zhou et al., 2023a), its utilization is obstructed by the aforementioned constraints. Hence, designing a FedRec that integrate multimodal data while balancing efficiency, privacy, heterogeneity, and sparsity is thus a critical research problem (Feng et al., 2024; Li et al., 2024a).
To address these interconnected challenges, we propose FedMR, a novel Federated Multimodal Recommendation framework. FedMR utilizes a collaborative server-client architecture where the server leverages powerful pre-trained models for centralized, efficient encoding of item modalities, creating rich, shared contextual representations that combat client-side item understanding limitations due to sparsity. Crucially, FedMR introduces the Mixing Feature Fusion Module (MFFM), operating on the client-side for dynamic and personalized adaptation. The MFFM empowers each client to learn user-specific weights for combining different fusion perspectives of the server-provided multimodal representations and collaborative signals from ID embeddings, guided by the client’s unique interaction history. This adaptive multimodal fusion directly tackles data heterogeneity by tailoring the final item representation to individual user preferences and local data characteristics, leading to more accurate and personalized recommendations while preserving privacy.
Our main contributions are summarized as follows:
  • Framework Innovation: We propose FedMR, a novel federated architecture employing centralized server-side processing of item modalities to enrich item context and alleviate client-side limitations while protecting user privacy.
  • Dynamic Feature Fusion: We introduce the Mixing Feature Fusion Module for dynamic, personalized adaptation of multimodal representations, effectively addressing data heterogeneity and diverse modality preferences.
  • Seamless Compatibility and Enhancement: FedMR seamlessly integrates with existing ID-based FedRec frameworks, upgrading them with multimodal capabilities while requiring minimal modification and boosting performance.
  • Comprehensive Empirical Validation: Extensive experiments on real-world multimodal datasets validate FedMR’s effectiveness, showing significant improvements over FedRec baselines. We also provide code for reproducibility.
@@ 2.1. Federated Recommendation Systems
FedRecs aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020). Significant progress has been made in FedRec across various tasks, including rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024b), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and cold-start scenarios (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d; Li et al., [n. d.]). However, many existing FedRec approaches (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) primarily use ID embeddings for item representations, thereby neglecting rich multimodal information like associated text and images. Recent efforts have started applying multimodal learning (Baltrušaitis et al., 2018) to FedRec, called Federated Multimodal Recommendation System (FedMMRec). AMMFRS (Feng et al., 2024) employs an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data based on feature importance. In contrast, FedMMR (Li et al., 2024a) offloads heavy multimodal computation to the server and employs secure cross-client convolution with secret sharing to capture high-order interactions privately. Other works focus on enhancing personalization within FedRec. PFedRec (Zhang et al., 2023c) introduced dual personalization for both users and items. FedRAP (Li et al., 2023a) refined this with additive personalization. FedDAE (Li et al., 2024c) uses a variational autoencoder and a gating network to adaptively fuse shared and personalized item knowledge. Furthermore, recent works explore integrating Foundation Models (FMs) directly within FedRec. FedPA (Zhang et al., 2024c) introduces personalized adapters for lightweight, client-specific FM fine-tuning. TransFR (Zhang et al., 2024b) uses FMs to generate universal textual item representations, replacing traditional IDs to improve adaptability and address cold-start issues. Despite these advancements in personalization and FM integration, methods predominantly relying on ID embeddings or single-modality representations still limit the full utilization of rich, multi-faceted multimodal item information within the federated setting (Che et al., 2023).
@@ 2.2. Multimodal Recommendation Systems
Traditional Recommendation Systems (RecSys) primarily rely on user-item interactions (Yu et al., 2023). In contrast, Multimodal Recommendation systems (MMRec) enhance accuracy and user experience by leveraging multiple data modalities like text, images, audio, and video (Zhou et al., 2023b). Item IDs, often used as unique identifiers, can capture collaborative signals and structural relationships (Yuan et al., 2023b; Liu et al., 2023). In MMRec, these IDs are frequently combined with multimodal data as auxiliary information to create richer item representations (Wang et al., 2024; Zhang et al., 2024e). This integration helps capture a more comprehensive understanding of users’ interests and preferences, ultimately improving recommendation quality. A typical MMRec process involves two main stages: feature extraction and modality fusion (Guo et al., 2019; Gao et al., 2020). Feature extraction (Guo et al., 2019) aims to obtain representative features from various modalities. Foundation Models (FMs), typically pre-trained on large-scale unlabeled datasets (Bommasani et al., 2021), have become increasingly prominent for this task. They can be adapted for recommendation through various techniques, including representation transfer (Du et al., 2020; Liu et al., 2021a, 2022; Wei et al., 2023), direct fine-tuning (Zhang et al., 2021; Li et al., 2022b, 2023b; Sun et al., 2023; Wang et al., 2023), parameter-efficient adapter tuning (Hou et al., 2022; Bian et al., 2023; Geng et al., 2023; Fu et al., 2024), or prompt-based tuning (Shin et al., 2020; Li and Liang, 2021; Zhou et al., 2022; Zhang and Wang, 2023; Xu et al., 2024). Several studies leverage FMs to enhance the content understanding capabilities of MMRec, often using cross-modal alignment or generation techniques. CLIP (Radford et al., 2021a; Wu et al., 2024), for example, uses contrastive learning (Le-Khac et al., 2020) between language and images to obtain powerful multimodal representations. M6-Rec (Cui et al., 2022) is an FM-based system tailored for e-commerce recommendations. P5 (Geng et al., 2022) frames recommendation as a natural language task solvable via prompts. VIP5 (Geng et al., 2023) utilizes pre-trained multimodal models to capture complementary visual and textual information for personalized recommendations. Following feature extraction, modality fusion (Gao et al., 2020) focuses on effectively integrating these features to enhance overall performance. Recommendation models then combine these fused multimodal features with interaction data to predict user preferences (Tao et al., 2020; Wu et al., 2021; Zhou and Shen, 2023). However, realizing the full potential of MMRec within the constraints of FL, especially regarding personalized and adaptive fusion across heterogeneous clients, necessitates specialized frameworks are needed to provide personalized recommendation with the privacy protection.
@@ 3. Problem Formulation
Let $\mathcal{U}$ and $\mathcal{I}$ denote the sets of users and items, respectively. User-item interactions are represented by a binary matrix $\mathbf{R}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|}$, where $r_{u,i}=1$ signifies an interaction between user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, and $r_{u,i}=0$ otherwise. The interaction history for user $u$ is denoted by the vector $\mathbf{r}_{u}\in\{0,1\}^{|\mathcal{I}|}$, and the set of observed interactions is $\mathcal{O}=\{(u,i)\mid r_{u,i}=1\}$. Each item $i\in\mathcal{I}$ is associated with features from a set of modalities $\mathcal{M}$. For the $m$-th modality, these features are represented by a matrix $\mathbf{E}_{m}\in\mathbb{R}^{|\mathcal{I}|\times d_{m}}$, where the vector $\mathbf{e}_{i,m}\in\mathbb{R}^{d_{m}}$ is the feature embedding for item $i$ in modality $m$. To handle multimodal information effectively, we introduce a set of modality fusion strategies $\mathcal{G}=\{g_{1},\dots,g_{|\mathcal{G}|}\}$ for dynamically combining features. We primarily focus on visual $v$ and textual $c$ modalities, i.e., =$\mathcal{M}=\{v,c\}$, although the FedMR framework is extensible to other modalities and fusion mechanisms.
@@ 4. Methodology
FedMR aims to seamlessly integrate multimodal item features with ID embeddings within a federated environment, enhancing existing ID-based FedRecs, to predict items that users have not yet explored but may find interesting. It centralizes multimodal data processing on the server for efficiency and rich item context, then uses dynamic client-side personalization via MFFM to handle heterogeneity.
4.1. Preliminaries
FedRec systems aim to collaboratively train recommendation models while preserving user privacy. The standard objective is to minimize the aggregated reconstruction loss across clients $u\in\mathcal{U}$:
(1) $\min\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(\hat{r}_{u,i},r_{u,i}),$
where $\mathcal{L}_{\text{recon}}$ is the loss function, and $\alpha_{u}$ is the weight of client $u$’s contribution. Many conventional FedRec methods (Lin et al., 2020a, b; Liang et al., 2021; Perifanis and Efraimidis, 2022; Zhang et al., 2023d) rely solely on a global item ID embedding matrix $\mathbf{D}$ and user-specific parameters $\theta_{u}$ to predict a score $\hat{r}_{ui}$ using a prediction function $f$:
(2) $\hat{r}_{ui}=f(\mathbf{D}_{i};\theta_{u}).$
Here, $\mathbf{D}_{i}$ represents the ID embedding for item $i$. These methods, so-called ID-based FedRecs, primarily capture collaborative signals and structural relationships (Yuan et al., 2023b), but often overlook rich multimodal item content, which motivating us to propose the FedMR.
4.2. Server-Side Multimodal Encoding
To efficiently incorporate multimodal item content while mitigating client resource limitations and communication overhead, FedMR performs feature encoding centrally on the server. Leveraging a pre-trained FM $h(\cdot)$, FedMR encodes raw visual features $\mathbf{E}_{v}$ and textual contents $\mathbf{E}_{c}$ into dense semantic embeddings: